# Drawback of Enforcing Equivariance and its Compensation via the Lens of Expressive Power

### v1

@@ 6.2 Non-realizable case: Enlarging model size can lead to even better expressive power
Let $s$ be any invariant target function, and $\mathcal{Q}f$ the projection of $f$ to the equivariant space that is defined as $\mathcal{Q}f=\frac{1}{|G|}\sum_{g\in G}f\circ\rho_{g}$. If the data distribution is symmetric, then we have that
@@
for all approximator $f$. Specifically, if $f$ is an output function of GNs, we can find that $\mathcal{Q}f$ can be an output function of GENs/LENs with a $|G|$ times model size, implying that GENs/LENs can fix the expressive power gap and achieve a smaller expected loss by a larger model size.
When the GNs are of the output function $f(x)=\sum_{i=1}^{n}\beta_{i}\sigma(\langle\alpha_{i},x\rangle)$, the weight matrices of GENs/LENs can be formulated as
$W^{(2)}=\frac{1}{|G|}[\underbrace{\beta_{1},\dots,\beta_{1}}_{|G|},\dots,\underbrace{\beta_{m},\dots,\beta_{m}}_{|G|}]~\text{and}$
@@
which implies a comparable expressive power but a $|G|$ times model size. Meanwhile, the weight matrices can be generated in two steps: (1) copy the same number of original channel vectors, and (2) generate all symmetric channel vectors to make the network equivariant. Therefore, although LENs have a larger model size, they have the same number of learnable parameters. As a result, the dimension of the hypothesis space can be decreased by choosing less original channel vectors or constrain the original channel vectors in a subspace, while ensuring a comparable expressive power.
If we use the method in Lemma 6.1 to compress the networks, we can further get a better result. More precisely, we can let the weight matrices be
$W^{(2)}=[\underbrace{\beta_{1},\dots,\beta_{1}}_{C_{1}},\dots,\underbrace{\beta_{m},\dots,\beta_{m}}_{C_{m}}]~\text{and}$
@@
where $\alpha_{i}$ is constrained to satisfy $|\{\rho_{g}^{T}\alpha_{i}:g\in G\}|=C_{i}$ and $g_{ij}$} is chosen to satisfy $\{\rho_{g_{i,j}}^{T}\alpha_{i}:j\in[C_{i}]\}=\{\rho_{g}^{T}\alpha_{i}:g\in G\}$. Similarly, this model still has comparable expressive power, but a lower-dimensional hypothesis space since we constrain all first-layer channel vectors in a (sub-)space.
Revise Example 5.1 where $C_{1}=4$ with arbitrary channel vector $(c_{1},c_{2})$ and $C_{2}=2$ with constrained channel vector $c_{3}(1,-1)$ that satisfies $\rho_{gh}(1,-1)=(1,-1)$. In this case, only two channel vectors are used originally to generate symmetric channel vectors, compared with $3$ for GNs. Moreover, the second channel vector $c_{3}(1,-1)$ are constrained in a one-dimension space, which further decreases the dimension of hypothesis space. As a result, the total numbers of parameters become
$\underbrace{2}_{\alpha_{1}}+\underbrace{1}_{\alpha_{2}}+\underbrace{2}_{\beta}\text{v.s.}\underbrace{2}_{\alpha_{1}}+\underbrace{2}_{\alpha_{2}}+\underbrace{2}_{\alpha_{3}}+\underbrace{3}_{\beta}.$
6.3 Constructed LEN with better generalizability
While increasing model size allows LENs to achieve comparable expressive power to GNs, a common concern is that this may harm generalization. In this section, we revisit Example 5.1 to demonstrate that despite requiring more neurons, LENs can still maintain a less complex hypothesis space than GNs. This lower complexity is significant as it suggests that LENs can achieve better generalization.
As discussed in Section 5, a LEN with 6 neurons and a GN with 3 neurons can both represent the target function. With the permutation representation $\psi$ defined as:
@@ 6.3 Constructed LEN with better generalizability
The output function of the LEN can be formulated as:
@@
From this formulation, we can determine that the dimension of the hypothesis space for the LEN (with 6 neurons) is 5. However, a GN with 3 neurons is defined as $\sum_{i=1}^{3}\beta_{i}\sigma(c_{1}x+c_{2}y)$, which means the 9-dimensional hypothesis space of the GN hypothesis space.
The key insight here is that although LENs may have a larger model size to maintain the expressive power, their inherent structural constraints could lead to a more lower-dimensional hypothesis space. Consequently, this reduced complexity implies that LENs are expected to exhibit superior generalization capabilities compared to standard GNs in this scenario.
@@ 7 Conclusions
In this work, we investigated the theoretical properties of equivariant neural networks, specifically focusing on the 2-layer ReLU architectures. Through the construction of an example, we demonstrated a key drawback of equivariance: it strictly reduces expressive power. Nevertheless, we proved that this reduction is recoverable by increasing the model size. We also show that although the larger model size is required to maintain expressivity, it could retain a lower complexity for the resulting hypothesis space, thereby still offering a superior generalizability for equivariant networks.
@@
Moreover, when $\psi_{ij}\not=0$, we can set $x=e_{j}$. Then $\sigma(\psi x)=\psi\sigma(x)$ implies that $\sigma(\psi_{ij})=\psi_{ij}$. From the definition of ReLU, we have that $\psi_{ij}>0$. The proof is completed.
@@
As a result, $M$ and $\rho_{g}M$ have the equivaENt conditions to be boundary hyperplanes, causing that $M$ is a boundary hyperplane if and only if $\rho_{g}M$ is a boundary hyperplane. The proof is completed.
@@
which is a satisfied rewritten LEN.
@@ Appendix E Proof of ENs’ Universal Represetation Ability
where the group representation $\psi$ in the intermediate layers is the corresponding permutation representation acting on each axis $e_{g_{i}}$ as $\psi_{g}e_{g_{i}}=e_{g_{i}g^{-1}}$. As a result, we have
@@
$\tilde{W}^{(\ell)}\psi_{g}=\text{diag}(\underbrace{W^{(\ell)},\dots,W^{(\ell)}}_{|G|})\psi_{g}=\text{diag}(\underbrace{W^{(\ell)},\dots,W^{(\ell)}}_{|G|})=\psi_{g}\text{diag}(\underbrace{W^{(\ell)},\dots,W^{(\ell)}}_{|G|})=\psi_{g}\tilde{W}^{(\ell)},\forall\ell=2,3\dots,L-1$
$\tilde{W}^{(L)}\psi_{g}=\frac{1}{|G|}(\underbrace{\phi_{g_{1}}^{-1}W^{(L)},\dots,\phi_{g_{|G|}}^{-1}W^{(L)}}_{|G|})\psi_{g}=\frac{1}{|G|}(\underbrace{\phi_{g_{1}g^{-1}}^{-1}W^{(L)},\dots,\phi_{g_{|G|g^{-1}}}^{-1}W^{(L)}}_{|G|})=\frac{1}{|G|}\phi_{g}(\underbrace{\phi_{g_{1}}^{-1}W^{(L)},\dots,\phi_{g_{|G|}}^{-1}W^{(L)}}_{|G|})$
As a result, the constructed network has the same output function $\mathcal{Q}F$ and equivariant layers, renders it an EN. The proof is completed.

### v2

@@ 6.2 Non-realizable Case
Let $s$ be any invariant target function. Denote by $\mathcal{Q}f$ the projection of $f$ to the equivariant space, defined as $\mathcal{Q}f=\frac{1}{|G|}\sum_{g\in G}f\circ\rho_{g}$. If the data distribution is symmetric, then we have that
@@
for all approximator $f$. Specifically, if $f$ is an output function of GNs, $\mathcal{Q}f$ can be an output function of GENs/LENs with a $|G|$ times model size, implying that GENs/LENs can fix the expressive power gap and achieve a smaller expected loss by a larger model size.
When the GN has an output function $f(x)=\sum_{i=1}^{n}\beta_{i}\sigma(\langle\alpha_{i},x\rangle)$, the weight matrices of GENs/LENs can be formulated as
$W^{(2)}=\frac{1}{|G|}[\underbrace{\beta_{1},\dots,\beta_{1}}_{|G|},\dots,\underbrace{\beta_{m},\dots,\beta_{m}}_{|G|}],\text{ and}$
@@
which implies comparable expressive power while requiring a model size proportional to $|G|$. Meanwhile, the weight matrices can be generated in two steps: (1) copy the same number of original channel vectors, and (2) generate all symmetric channel vectors to make the network equivariant. Therefore, although LENs have a larger model size, they have the same number of learnable parameters. As a result, the dimensionality of the hypothesis space can be reduced by selecting fewer original channel vectors or by constraining them to a subspace, while maintaining comparable expressive power.
If we use the method in Lemma 6.1 to compress the networks, we can further get an even better result. More precisely, we can let the weight matrices be
$W^{(2)}=[\underbrace{\beta_{1},\dots,\beta_{1}}_{C_{1}},\dots,\underbrace{\beta_{m},\dots,\beta_{m}}_{C_{m}}],\text{ and}$
@@
where $\alpha_{i}$ is constrained to satisfy $|\{\rho_{g}^{T}\alpha_{i}:g\in G\}|=C_{i}$, and $g_{ij}$} is chosen to satisfy $\{\rho_{g_{i,j}}^{T}\alpha_{i}:j\in[C_{i}]\}=\{\rho_{g}^{T}\alpha_{i}:g\in G\}$. Similarly, this model still has comparable expressive power, but a lower-dimensional hypothesis space, since we constrain all first-layer channel vectors within a (sub-)space.
6.3 Enlarged LEN with Smaller Hypothesis Complexity
While increasing model size allows LENs to achieve comparable expressive power to GNs, a natural concern is that this may harm the generalizability. In this section, we revisit Example 5.1 to demonstrate that despite requiring more neurons, LENs can still maintain a less complex hypothesis space than GNs. This lower complexity is significant as it suggests that LENs can achieve better generalizability.
As discussed in Section 5, a LEN with six neurons and a GN with three neurons can represent the same target function. With the permutation representation $\psi$ defined as
@@ 6.3 Enlarged LEN with Smaller Hypothesis Complexity
Then the output function of the LEN can be formulated as
@@
From this formulation, we can determine that is characterized by five parameters, meaning its parameter space is $\mathbb{R}^{5}$. In contrast, a three-neuron GN is defined as $\sum_{i=1}^{3}\beta_{i}\sigma(c_{1}x+c_{2}y)$, which involves 9 parameters, resulting in a larger parameter space of $\mathbb{R}^{9}$.
The key insight here is that although LENs may have a larger model size to maintain the expressive power, their inherent structural constraints could lead to a lower-dimensional hypothesis space. Consequently, this reduced complexity implies that LENs are expected to exhibit superior generalizability compared to standard GNs in this scenario.
@@ 7 Conclusions
Equivariant neural networks leverage data symmetries as a structural inductive bias, leading to significant success across various applications. Despite their popularity, their expressive power is not yet fully understood. This study evaluates how enforcing equivariance influences the expressive power of two-layer ReLU networks. By analyzing boundary hyperplanes and channel vectors, we show that these constraints can inherently limit a model’s expressive power. We demonstrate, however, that this limitation can be compensated for by increasing model scale, through formal upper bounds for the necessary expansion. Notably, our findings reveal that these larger equivariant architectures possess a reduced hypothesis space dimensionality, suggesting superior generalization potential.
@@
Moreover, when $\psi_{ij}\not=0$, we can set $x=e_{j}$. Then $\sigma(\psi x)=\psi\sigma(x)$ implies that $\sigma(\psi_{ij})=\psi_{ij}$. From the definition of ReLU, we have that $\psi_{ij}>0$. The proof is completed. ∎
@@
As a result, $M$ and $\rho_{g}M$ have the equivaENt conditions to be boundary hyperplanes, causing that $M$ is a boundary hyperplane if and only if $\rho_{g}M$ is a boundary hyperplane. The proof is completed. ∎
@@
which is a satisfied rewritten LEN. ∎
@@ Appendix E Proof of ENs’ Universal Represetation Ability
where the group representation $\psi$ in the intermediate layers is the corresponding permutation representation acting on each axis $e_{g_{i}}$ as $\psi_{g}e_{g_{i}}=e_{g_{i}g^{-1}}$. As a result, we have, $\forall\ell=2,3\dots,L-1$,
@@
$\tilde{W}^{(\ell)}\psi_{g}=$ $\text{diag}(\underbrace{W^{(\ell)},\dots,W^{(\ell)}}_{|G|})\psi_{g}=\text{diag}(\underbrace{W^{(\ell)},\dots,W^{(\ell)}}_{|G|})=\psi_{g}\text{diag}(\underbrace{W^{(\ell)},\dots,W^{(\ell)}}_{|G|})=\psi_{g}\tilde{W}^{(\ell)}$,
$\tilde{W}^{(L)}\psi_{g}=$ $\frac{1}{|G|}(\underbrace{\phi_{g_{1}}^{-1}W^{(L)},\dots,\phi_{g_{|G|}}^{-1}W^{(L)}}_{|G|})\psi_{g}=\frac{1}{|G|}(\underbrace{\phi_{g_{1}g^{-1}}^{-1}W^{(L)},\dots,\phi_{g_{|G|g^{-1}}}^{-1}W^{(L)}}_{|G|})$
$=$ $\frac{1}{|G|}\phi_{g}(\underbrace{\phi_{g_{1}}^{-1}W^{(L)},\dots,\phi_{g_{|G|}}^{-1}W^{(L)}}_{|G|})$
As a result, the constructed network has the same output function $\mathcal{Q}F$ and equivariant layers, renders it an EN. The proof is completed. ∎